OpenStudy (anonymous):

Together a computer, monitor, and printer cost $1,935. The monitor cost $400 less than the computer, and the printer cost $50 more than the monitor. Find the cost of the printer. ~Equation

OpenStudy (anonymous):

x is the computer, x-400 is the monitor x-400+50 is the cost of printer

OpenStudy (anonymous):

Need more help?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

ok x+x-400 + x-400+50 is the total cost of all three so 3x-750 = 1935 3x = 2685 x = 895 now you have the cost all three so $895 cost of computer so printer was x-400+50 that is 895-400+50 =?

OpenStudy (anonymous):

545

OpenStudy (anonymous):

or 445

OpenStudy (anonymous):

545 you were right the first time

OpenStudy (anonymous):

ok thats the price of?

OpenStudy (anonymous):

the printer

OpenStudy (anonymous):

i thought it was the moniter because the printer was 895 and the monitor is 545

OpenStudy (anonymous):

Ill type it out so it looks more reasonable for you

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

12 mins?

OpenStudy (anonymous):

i g2g its lie midnight

OpenStudy (anonymous):

bye

OpenStudy (anonymous):

post it and ill get to it in the morning

OpenStudy (anonymous):

Ok we are given that 1935 is the cost of EVERYTHING, so the computer + monitor + printer = $1935. Then they tell us that the monitor costs $400 LESS THAN the computer, now heres the trick with ALL these types of questions. We dont know how much the computer costs we are just told that the monitor is 400 LESS than it, so we give the computer some letter to represent what we dont know yet, say 'x' it could be anything you like though. So again we are told the monitor is 400 less than 'x' x being the cost of the computer, so we have monitor = x-400 dollars So far we still have no clue at all how much the computer is or the monitor, just that the monitor is x-400. Next we are told that the printer costs 50 dollars more than the monitor, ok so if the monitor is x-400 and the printer is 50 dollars MORE than the monitor that makes the printer worth monitor+50, so we have printer = monitor + 50 = x-400+50 ok so now we have values for the computer monitor and printer, yet we still have no idea how much they actually cost! However, they told us at the start how much it all cost, that is 1935, so we write it all down, that is monitor + printer + computer = 1935 so getting all our totals we have now monitor = x-400 printer = x-400+50 and the the computer was worth 'x' so we have x-400+x-400+50+x = 1935 now we want to simplify this a little, we have 3 'x's 3x-400-400+50 = 1935 3x-800+50 = 1935 3x-750 = 1935 ok we add 750 to both sides 3x = 2685 now divide both sides by 3 to get x by itself x = 2685/3 x = 895 Now we are in business, we have a value for the unknown 'x' remember we used 'x' as our value for the computer! Now we are in a position to calculate the price of ANY of the items. If we go back and look at our equations we see the monitor costs 400 less than the computer, we know that the computer now costs 895 so 400 less than that will be 895 - 400 = 495 So the monitor costs 495. Next we were told that the printer was 50 more than the monitor so we have 495+50 = 545

OpenStudy (anonymous):

Thx